# Learning Hierarchical Sparse Transform Coding of 3DGS

### v2

@@ 14 Additional Comparisons with Other Methods
In addition to the comparisons with anchor-based compression methods using R–D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines. These include several commonly used 3DGS and compression baselines [kerbl20233d, lu2024scaffold, lee2024compact, niedermayr2024compressed, girish2024eagles, fan2024lightgaussian, morgenstern2024compact, navaneet2024compgs, papantonakis2024reducing, wang2024end, liu2024compgs], the feed-forward compression method FCGS [chen2025fast], and two neural-field-based 3DGS compression methods, LocoGS [shin2025localityaware] and NeuralGS [tang2025neuralgs]. We further consider several recent 3DGS compression methods [hanson2025pup, xie2025sizegs, tian2025flexgaussian, lee2025optimized], two HAC-based approaches that introduce prediction modules to improve HAC [tang2025feature, ma2025enhancing], and a group of methods that apply transform coding in a post-training stage for 3DGS compression [morgenstern2024compact, lee2025compression, xie2024mesongs, yang2025hybridgs].
@@
Regarding SOGS [zhang2025sogs], we do not include it in our quantitative comparison for two reasons. First, to the best of our knowledge, there is no official implementation available, and the original paper does not report file sizes or distortion metrics on the datasets used in our experiments, which makes a fair and reproducible comparison difficult. Second, SOGS is essentially a 3DGS compaction method rather than a full compression scheme with an explicit bitstream and entropy coding. Based on the model statistics reported in [zhang2025sogs], we can only roughly infer an effective compression ratio of about $2\times$ over Scaffold-GS; this is merely a rough estimate, since the authors do not provide explicit compression ratios. In contrast, by employing the proposed SHTC framework, our method achieves substantially higher compression ratios over Scaffold-GS. For these reasons, we omit SOGS from the quantitative comparison.
@@
Because many of the above methods report only one or two operating points on the R–D curve, a meaningful BD-rate computation is not feasible. For these methods, we therefore provide a numerical comparison in tabular form in Tab. 7. As shown in Tab. 7, our method achieves a better memory-quality trade-off than above approaches.
@@
15 Implementation Details
15.1 Pixel-wise Error in YCbCr Color Space
@@
$\mathcal{L}_{\mathrm{YCbCr}}=$ $\ell_{1}(I_{Y},\hat{I}_{Y})+\lambda_{c}\ell_{1}(I_{Cb},\hat{I}_{Cb})+\lambda_{c}\ell_{1}(I_{Cr},\hat{I}_{Cr})$
$+$ $\lambda_{H}\ell_{1}(\nabla^{2}I_{Y},\nabla^{2}\hat{I}_{Y})$ (11)
$+$ $\lambda_{TV}\mathrm{TV}(\hat{I}_{Cb})+\lambda_{TV}\mathrm{TV}(\hat{I}_{Cr}).$
@@
15.2 Details of $\lambda$ Selection for R-D Curves
In this subsection, we summarize in Tab. 8 the $\lambda$ values used to generate the R–D curves in our experiments. In addition to these values, we also clarify several special cases regarding how $\lambda$ and the corresponding R-D points are obtained for some baselines:
  • Main text. For the R-D curves in the main paper, we train the model with $\lambda\in\{0.002,0.004,0.008,0.015\}$, which yields a PSNR range of roughly 1 dB. The spacing between consecutive $\lambda$ value is gradually increased so as to cover a wide bitrate range while providing approximately uniform sampling of the R–D curves within that range. Because different models and datasets exhibit different sensitivity to $\lambda$, we additionally use larger $\lambda$ values for some baselines so that the PSNR range of each curve is comparable to that of the others.
  • CAT-3DGS. For CAT-3DGS [zhan2025catdgs], we encountered out-of-memory (OOM) errors during rendering when using the official implementation on our hardware. Consequently, we do not re-train or re-evaluate this method; instead, we directly use the R-D points reported in the original paper as its results in our plots.
@@ 15.2 Details of $\lambda$ Selection for R-D Curves
  • PCGS. For PCGS, on both Synthetic-NeRF and BungeeNeRF, we found it difficult to obtain a sufficiently wide PSNR range for a stable BD-rate estimation by tuning $\lambda$. Therefore, we only plot the R-D points reported in the original paper and treat PCGS as a qualitative reference rather than including it in the quantitative BD-rate comparison.
16 Limitation and Future Work
@@
17 Supplementary Related Work
17.1 Unstructured 3DGS compression methods
@@
The densification process of 3DGS results in an explosion in the number of Gaussians. While this allows for the reconstruction of finer scene details, a substantial portion of the Gaussians is redundant, and removing them has little effect on visual quality. The challenge lies in identifying these unimportant Gaussians. [navaneet2024compgs] simply uses opacity as a criterion to remove transparent or nearly invisible Gaussians. [ali2024elmgs, ali2024trimming] further use both gradient magnitude and opacity levels as reference to identify removable Gaussians. Several subsequent studies assess each Gaussian’s contribution to ray color prediction and utilize this as the importance score [fan2024lightgaussian, girish2024eagles, niemeyer2024radsplat]. Lee et al. learn a binary mask to prune those unimportant Gaussians [lee2024compact], which is adopted by several subsequent works [wang2024end]. [hanson2025pup] propose a principled sensitivity pruning score by computing a second-order approximation of the reconstruction error with respect to the parameters of each Gaussian.
@@
The largest portion of the memory footprint is used to store SH coefficients for modeling view-dependent color. However, a large fraction of the region consists of diffuse materials, which can be effectively modeled with view-independent color. This leads to significant waste in using high-degree SHs. To address this, [papantonakis2024reducing] assign each Gaussian an appropriate SH degree to reduce waste, while [fan2024lightgaussian] distill knowledge from a high-degree SH teacher model to a student model with truncated, lower-degree SHs.
@@
Gaussian attributes are typically stored at 32-bit precision. However, opacity, scale, rotation, and SH coefficients are tolerant to small inaccuracies, suggesting that lower bit-depths can represent them with minimal performance drop. In light of this, [papantonakis2024reducing] propose a scalar quantization scheme that assigns each full-precision value to its nearest codeword in a K-means-derived codebook, storing only the corresponding index. [girish2024eagles] learn a latent representation, quantize it using a uniform scalar quantizer, and employ a decoder to predict Gaussian attributes from the quantized representation.
@@
Many Gaussians share similar parameters, allowing similar vectors to be mapped to a common code vector in a codebook. Instead of storing high-precision vectors, only compact indices are needed. Based on this insight, several works apply vector quantization to eliminate the redundancy within Gaussian attributes [navaneet2024compgs, fan2024lightgaussian, niedermayr2024compressed]. Furthermore, [wang2024end] employ entropy-constrained vector quantization, incorporating entropy coding into the quantization process. It assigns shorter codes to those code vectors that are used more frequently, effectively compressing the index stream.
17.2 Transform coding in NeRF/3DGS compression
Before the emergence of 3DGS, grid-based NeRF representations [muller2022instant, fridovich2023k, sun2022direct, chen2022tensorf] were considered a promising alternative to the classic NeRF [mildenhall2021nerf]. Several works have explored the use of transform coding to compress the feature grids in these representations [rho2023masked, li2024nerfcodec, pham2024neural]. Specifically, Rho et al. use a Discrete Wavelet Transform (DWT) to exploit transform-domain sparsity for compressing feature planes [rho2023masked]. Li et al. repurpose neural image codecs as learned nonlinear transforms: for each scene, they pre-train a plane-based NeRF, then scene-specifically fine-tune a subset of neural codec layers and encode the pretrained feature planes with the fine-tuned codec [li2024nerfcodec]. Pham et al. follow the same two-stage recipe but replace the off-the-shelf codec with lightweight nonlinear transforms designed from scratch [pham2024neural].
Compared with grid-based NeRFs, applying transform coding to 3DGS is more challenging because 3DGS data are irregular and unordered. To sidestep these challenges, prior work defers the application of transform coding until the post-training stage, once the 3DGS geometry is fixed. Two main strategies have been explored. First, some methods project 3DGS onto a set of feature planes [morgenstern2024compact] or adopt a triplane representation [chan2022efficient] to implicitly encode Gaussian attributes [lee2025compression], then compress the learned planes with standard image or video codecs. Second, other methods treat a pretrained 3DGS model as a point cloud and apply point-cloud attribute transforms such as Graph Fourier transform (GFT) [sandryhaila2013discrete] and Region-adaptive Hierarchical Transform (RAHT) [de2016compression] to compress the model [xie2024mesongs, yang2024benchmark, huang2025hierarchical, wang2025adaptive].
The above methods borrow transforms from images, video, or point clouds but apply fixed, linear transforms only after training. This decoupling blocks mutual adaptation between the transform and the 3DGS representation, often limiting R–D performance. In contrast, we propose an end-to-end transform-coding framework that jointly adapts the transform and the 3DGS representation, improving compression performance over post-training transform approaches.

### v3

@@ Appendix D Additional Comparisons with Other Methods
In addition to the comparisons with anchor-based compression methods using R–D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines. These include several commonly used 3DGS and compression baselines (Kerbl et al., 2023; Lu et al., 2024; Lee et al., 2024; Niedermayr et al., 2024; Girish et al., 2024; Fan et al., 2024; Morgenstern et al., 2024; Navaneet et al., 2024; Papantonakis et al., 2024; Wang et al., 2024a; Liu et al., 2024b), the feed-forward compression method FCGS (Chen et al., 2025b), and two neural-field-based 3DGS compression methods, LocoGS (Shin et al., 2025) and NeuralGS (Tang et al., 2025b). We further consider several recent 3DGS compression methods (Hanson et al., 2025; Xie et al., 2025; Tian et al., 2025; Lee et al., 2025a), two HAC-based approaches that introduce prediction modules to improve HAC (Tang et al., 2025a; Ma et al., 2025), and a group of methods that apply transform coding in a post-training stage for 3DGS compression (Morgenstern et al., 2024; Lee et al., 2025b; Xie et al., 2024; Yang et al., 2025).
@@
Regarding SOGS (Zhang et al., 2025), we do not include it in our quantitative comparison for two reasons. First, to the best of our knowledge, there is no official implementation available, and the original paper does not report file sizes or distortion metrics on the datasets used in our experiments, which makes a fair and reproducible comparison difficult. Second, SOGS is essentially a 3DGS compaction method rather than a full compression scheme with an explicit bitstream and entropy coding. Based on the model statistics reported in (Zhang et al., 2025), we can only roughly infer an effective compression ratio of about $2\times$ over Scaffold-GS; this is merely a rough estimate, since the authors do not provide explicit compression ratios. In contrast, by employing the proposed SHTC framework, our method achieves substantially higher compression ratios over Scaffold-GS. For these reasons, we omit SOGS from the quantitative comparison.
@@
Because many of the above methods report only one or two operating points on the R–D curve, a meaningful BD-rate computation is not feasible. For these methods, we therefore provide a numerical comparison in tabular form in Table 4. As shown in Table 4, our method achieves a better memory-quality trade-off than above approaches.
[…]
@@
Appendix E Implementation Details
The overall training loss is defined as
[…]
E.1 Pixel-wise Error in YCbCr Color Space
@@
$\mathcal{L}_{\mathrm{YCbCr}}=\ell_{1}(I_{Y},\hat{I}_{Y})+\lambda_{c}\ell_{1}(I_{Cb},\hat{I}_{Cb})+\lambda_{c}\ell_{1}(I_{Cr},\hat{I}_{Cr})+\lambda_{H}\ell_{1}(\nabla^{2}I_{Y},\nabla^{2}\hat{I}_{Y})+\lambda_{TV}\mathrm{TV}(\hat{I}_{Cb})+\lambda_{TV}\mathrm{TV}(\hat{I}_{Cr})$ (10)
@@
E.2 Details of $\lambda$ Selection for R-D Curves
In this subsection, we summarize the $\lambda$ values used to obtain the R-D curves in our experiments in Table 5. For our primary experiments, we set the Lagrange multiplier $\lambda$ to $\{0.002,0.004,0.008,0.015\}$, which yields operating points spanning a wide rate-distortion range and enables a comprehensive comparison. To obtain more uniformly spaced operating points along the R–D curve, we use progressively larger gaps between successive $\lambda$ values. This is motivated by the highly nonlinear relationship between $\lambda$ and the achieved rate. In particular, in the low-rate regime, larger changes in $\lambda$ are often necessary to produce clearly separated rates. In addition, we discuss several special cases for baselines, detailing how $\lambda$ (or its equivalent control parameter) and the corresponding R–D points are obtained.
  • CAT-3DGS. For CAT-3DGS (Zhan et al., 2025), we encountered out-of-memory (OOM) errors during rendering when using the official implementation on our hardware. Consequently, we do not re-train or re-evaluate this method; instead, we directly use the R-D points reported in the original paper as its results in our plots.
@@ E.2 Details of $\lambda$ Selection for R-D Curves
  • PCGS. For PCGS (Chen et al., 2025a), on DeepBlending, Mip-NeRF360, and Tanks&Temples, we train the model with $\lambda\in\{0.004,0.008,0.015,0.025\}$ to enlarge the distortion range. Nevertheless, even with this wider distortion span, we still cannot obtain sufficient PSNR overlap with other methods, and we find it generally difficult to further increase the overlap for a meaningful BD-rate computation. Therefore, we include these PCGS results only for plotting the R-D curves as a qualitative comparison, and we do not include PCGS in the quantitative BD-rate evaluation. Given that expanding the overlap is consistently challenging on the above three datasets, for Synthetic-NeRF and BungeeNeRF we do not introduce new $\lambda$ values for re-training; instead, we directly use the results reported by the original authors as a reference in the R-D plots (also excluded from BD-rate). We note that our evaluation setup is largely comparable to PCGS, since both are derived from the HAC/HAC++ codebase.
Appendix F Limitation and Future Work
@@
Appendix G Supplementary Related Work
G.1 Unstructured Compression Paradigm
@@
The densification process of 3DGS results in an explosion in the number of Gaussians. While this allows for the reconstruction of finer scene details, a substantial portion of the Gaussians is redundant, and removing them has little effect on visual quality. The challenge lies in identifying these unimportant Gaussians. (Navaneet et al., 2024) simply uses opacity as a criterion to remove transparent or nearly invisible Gaussians. (Ali et al., 2024a, b) further use both gradient magnitude and opacity levels as reference to identify removable Gaussians. Several subsequent studies assess each Gaussian’s contribution to ray color prediction and utilize this as the importance score (Fan et al., 2024; Girish et al., 2024; Niemeyer et al., 2024). Lee et al. learn a binary mask to prune those unimportant Gaussians (Lee et al., 2024), which is adopted by several subsequent works (Wang et al., 2024a). (Hanson et al., 2025) propose a principled sensitivity pruning score by computing a second-order approximation of the reconstruction error with respect to the parameters of each Gaussian.
@@
The largest portion of the memory footprint is used to store SH coefficients for modeling view-dependent color. However, a large fraction of the region consists of diffuse materials, which can be effectively modeled with view-independent color. This leads to significant waste in using high-degree SHs. To address this, (Papantonakis et al., 2024) assign each Gaussian an appropriate SH degree to reduce waste, while (Fan et al., 2024) distill knowledge from a high-degree SH teacher model to a student model with truncated, lower-degree SHs.
@@
Gaussian attributes are typically stored at 32-bit precision. However, opacity, scale, rotation, and SH coefficients are tolerant to small inaccuracies, suggesting that lower bit-depths can represent them with minimal performance drop. In light of this, (Papantonakis et al., 2024) propose a scalar quantization scheme that assigns each full-precision value to its nearest codeword in a K-means-derived codebook, storing only the corresponding index. (Girish et al., 2024) learn a latent representation, quantize it using a uniform scalar quantizer, and employ a decoder to predict Gaussian attributes from the quantized representation.
@@
Many Gaussians share similar parameters, allowing similar vectors to be mapped to a common code vector in a codebook. Instead of storing high-precision vectors, only compact indices are needed. Based on this insight, several works apply vector quantization to eliminate the redundancy within Gaussian attributes (Navaneet et al., 2024; Fan et al., 2024; Niedermayr et al., 2024). Furthermore, (Wang et al., 2024a) employ entropy-constrained vector quantization, incorporating entropy coding into the quantization process. It assigns shorter codes to those code vectors that are used more frequently, effectively compressing the index stream.
G.2 Post-training Transform Coding Paradigm
In this subsection, we focus on methods that employ transforms during the post-training stage of 3DGS compression.
Applying transform coding to 3DGS is challenging because 3DGS data are irregular and unordered. To sidestep these challenges, prior work defers the application of transform coding until the post-training stage, once the 3DGS geometry is fixed. Two main strategies have been explored. First, some methods project 3DGS onto a set of feature planes or adopt a triplane representation (Chan et al., 2022) to implicitly encode Gaussian attributes, then compress the learned planes with standard image or video codecs (Lee et al., 2025b; Morgenstern et al., 2024). Second, other methods treat a pretrained 3DGS model as a point cloud and apply point-cloud attribute transforms such as Graph Fourier transform (GFT) (Sandryhaila and Moura, 2013) and Region-adaptive Hierarchical Transform (RAHT) (De Queiroz and Chou, 2016) to compress the model (Xie et al., 2024; Yang et al., 2024; Huang et al., 2025; Wang et al., 2025; Yang et al., 2025). The above methods borrow transforms from images, video, or point clouds but apply fixed, linear transforms only after training. This decoupling blocks mutual adaptation between the transform and the 3DGS representation, often limiting R–D performance. In contrast, we propose an end-to-end transform-coding framework that jointly adapts the transform, entropy model and the 3DGS representation, significantly improving compression performance over post-training transform approaches.